40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9157     Entry time: Tue Sep 24 22:19:57 2013
Author: Manasa 
Type: Update 
Category: General 
Subject: FPMI locked 

 [Masayuki, Manasa]

We locked FPMI and measured the FPMI noise (power spectrum of error signal - MICH_IN1) which will be calibrated.

The arms were locked using POX11 and POY11. The sign of MICH gain was changed to lock FPMI (from -30 to +30). 

ELOG V3.1.3-